We're upgrading Dragen to CentOS7 and installing the new filesystem. We will update the log when its back online.
The maintenance is starting at 14.30 and will last no more than 15 minutes.
There is an issue with password and QR code reset which we are fixing now.
The colossus file system is being upgraded from 10 - 24 June. During this time, no jobs can be run.
We are debugging and fixing the issue.
Colossus storage is currently down. We are addressing the issue, working to get jobs running asap.
TSD is having a network maintenance from 07:00 - 09:00 CET, and there will be interruptions to services during this period.
Colossus is currently down due to a crash in the cluster file system. We are working to resolve the issue. Submit nodes are also down as a side effect of this ongoing issue.
The PyPI, CRAN, STATA etc mirrors are down. We working on bringing them back online.
Some compute nodes are down at the moment, causing jobs to be re-queued. We are working to bring them back online.
There are currently problems with the file-export feature on data.tsd.usit.no
The issue should now be resolved.
Currently, some users are not able to log in to TSD via ThinLinc. We are working on fix.
There's a license upgrade issue.
Update: the new licenses have been installed.
The machine exporting the /cluster filesystem crashed, and was down for around 10 minutes. This has caused a lot of machines mounting this share to be unresponsive.
All machines which are having trouble mounting the share will be rebooted shortly to fix any issues.
All machines should now be available again. Let us know if you still have trouble accessing your machines.
Due to an issue with the virtualisation infrastructure, some VMs have been unintentionally rebooted. If you are having issues please let us know. We are investigating to fix any outstanding problems.
We need to perform a brief maintenance on the data portal, during which it will not be possible to import or export files.
We are investigating issues related to export in TSD projects
Due to a Linux operating system update some virtual machines are experiencing issues. We are working to solve the problem.
While this is ongoing, it will not be possible to reset passwords, or get new QR codes.
Since Saturday evening, there has been some instability with the Colossus file system, which affects all submit hosts and running jobs on Colossus.
The file system was up again at 09.44. All systems should work as normal, but please inform us if you have problems.
As a step in the process of getting the new storage in production, we will restrict access to Colossus from Monday March 23rd. to March 27th.
During this time, there will be no access to Colossus, and the /cluster storage, including the project areas and software modules. You can also not access these services from the TSD's virtual machines.
The HNAS areas ("durable") will not be affected by this downtime.
We will update this page with the progress of our work during the maintenance window.
Bio-IT processor software upgraded to v3.5.7.